Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

George is aiming his arrows at the target shown - probabilit

Expert replies
by GMIHIR » Sat May 19, 2012 8:16 am
George is aiming his arrows at the target shown by his mentor.If probability of hitting the target is 0.7, what is the approximate probability that he will hit the target only three times in his first five attempts?
A 0.1

B 0.3

C 0.85

D 0.5

E 0.65
Join the discussion
Source: — Problem Solving |

by GmatKiss » Sat May 19, 2012 9:06 am
GMIHIR wrote:George is aiming his arrows at the target shown by his mentor.If probability of hitting the target is 0.7, what is the approximate probability that he will hit the target only three times in his first five attempts?
A 0.1

B 0.3

C 0.85

D 0.5

E 0.65
Between C and E, E with a wild guess!
Join the discussion

by ayush.nitw » Sat May 19, 2012 10:32 am
Is the answer B?
as Hit = .7 & miss = .3
Binomial probability distribution = 5C3 (.7)^3 * (.3)^2
=.3087
Join the discussion

by Jeff@TargetTestPrep » Thu Dec 14, 2017 7:29 am
GMIHIR wrote:George is aiming his arrows at the target shown by his mentor.If probability of hitting the target is 0.7, what is the approximate probability that he will hit the target only three times in his first five attempts?
A 0.1

B 0.3

C 0.85

D 0.5

E 0.65
We need to determine the probability that George hits the target 3 times and misses twice, in 5 attempts.

Since the probability of a hit is 0.7, the probability of a miss is 0.3. Thus, we need to determine:

P(H-H-H-M-M) = 0.7 x 0.7 x 0.7 x 0.3 x 0.3 ≈ 0.03

However, we have to account for the number of ways we can arrange H-H-H-M-M, which is a combination problem because order doesn’t matter.

5C3 = 5!/[3!(5-3)!] = 5!/(3! x 2!) = (5 x 4)/2 = 10 ways

Thus, P(3H and 2M) ≈ 0.03 x 10 = 0.3.

Answer: B

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion